### v1

# Manifolds with harmonic curvature and curvature operator of the second kind

Abstract.
We prove that compact Riemannian manifolds of dimension $n\geq 3$ with harmonic curvature and $\frac{n(n+2)}{2(n+1)}$-nonnegative curvature operator of the second kind must be Einstein. In particular, Building upon Dai-Fu’s work [DF], it follows that if the curvature operator of the second kind is $\min\{\frac{n(n+2)}{2(n+1)},\max\{4,\frac{(n+2)}{4}\}\}$-nonnegative , then such a manifold must be of constant curvature.
Supported in part by National Natural Science Foundations of China #12461008 and 12271069, Jiangxi Province Natural Science Foundation of China #20202ACB201001, Jiangxi Province Graduate Student Innovation Special Fund Project #YC2025-B037.
1. introduction
It is an important topic in geometry to understand how geometric assumptions restrict the topology of the underlying Riemannian manifold. In this direction, Tachibana [Ta] showed that any compact Riemannian manifold of dimension $n\geq 3$ with harmonic curvature and positive curvature operator is isometric to a quotient of the standard sphere. If the curvature operator is only nonnegative, then the manifold is locally symmetric. Later, Tran [tran] proved that a compact Riemannian manifold $M$ of dimension $n\geq 4$ with harmonic Weyl curvature and positive curvature operator, then $M$ is locally conformally flat. In [PW1] Petersen and Wink showed that a compact manifold with harmonic Weyl tensor and $\frac{n-1}{2}]-$nonnegative curvature operator is either globally conformal to a quotient of the standard sphere or locally symmetric, and that the first possibility always occurs if the curvature operator is $\frac{n-1}{2}]-$positive. In [PW] they proved a similar Tachibana-type theorem under the stronger condition that the manifold be Einstein. More recently, Colombo, Mariani, and Rigoli [CMR] established that a compact Riemannian manifold of dimension $n\geq 3$ with harmonic curvature and $\frac{n-1}{2}]-$positive curvature operator must have constant sectional curvature, thereby generalizing Tachibana’s result.
For curvature operator of the second kind, Nishikawa [Ni] conjectured that a closed simply connected Riemannian manifold for which the curvature operator of the second kind is nonnegative and positive is diffeomorphic to a Riemannian locally symmetric space and a spherical space form, respectively. This conjecture has been proved and improved by Cao–Gursky–Tran [CGT] and Li [Li4], who relaxed the assumption to 33-positive (33-nonnegative) curvature operator of the second kind. In 1993, Kashiwada [ka] showed that Riemannian manifolds with harmonic curvature for which the curvature operator of the second kind is nonnegative (respectively, positive) are locally symmetric spaces (respectively, constant curvature spaces). Nienhaus, Petersen and Wink [NP, NPWW] used the Bochner formula to show that $n$-dimensional compact Einstein manifolds with $k(<\frac{3n(n+2)}{2(n+4)})$-nonnegative curvature operators of the second kind are either rational homology spheres or flat. Dai-Fu [DF] proved that any closed Einstein manifold of dimension $n=4$ or $n=5$ or $n\geq 8$ with $\frac{n+2}{4}]-$nonnegative curvature operator of the second kind is a constant curvature space. Later, Dai-Fu-Yang [DFD] proved a compact manifold with harmonic Weyl tensor and nonnegative curvature operator of the second kind is globally conformally equivalent to a space of positive constant curvature or is isometric to a flat manifold.
Let $\lambda_{i}$ be the eigenvalues of the curvature operator, arranged in non-decreasing order. If
${\lambda}_{1}+{\lambda}_{2}+\cdots{\lambda}_{[k]}+(k-[k])\lambda_{[k]+1}>0(\geq 0),$
we say that the curvature operator is $k$-positive(resp. nonnegative). In this paper, we investigate Riemannian manifolds of dimension $n\geq 3$ with harmonic curvature and $\frac{n(n+2)}{2(n+1)}$-nonnegative curvature operator of the second kind, and obtain the following Theorem $1.$
Theorem $1.$
Let $(M,g)$ be an $n(\geq 3)$-dimensional compact Riemannian manifold with harmonic curvature. If the curvature operator of the second kind $\mathring{R}$ is $\frac{n(n+2)}{2(n+1)}$-nonnegative, then $M$ is an Einstein manifold.
Furthermore, by employing rigidity theorems for Einstein manifolds concerning the curvature operator of the second kind, established by Dai–Fu [DF] and Li [Li3], we get the following:
Corollary 1.2.
Let $(M,g)$ be an $n(\geq 3)$-dimensional compact Riemannian manifold with harmonic curvature, and $\mathring{R}\,$ be the curvature operator of the second kind. If
(i) $3\leq n\leq 7$ and $\mathring{R}\,$ is $\frac{n(n+2)}{2(n+1)}$-nonnegative;
(ii) $8\leq n\leq 13$ and $\mathring{R}\,$ is 44-nonnegative;
(iii) $n\geq 14$, $\mathring{R}\,$ is $\frac{n+2}{4}]-$nonnegative,
then $M$ must be of constant curvature.
Remark 1.3.
Corollary 1.2 extends and enhances Kashiwada’s results [ka] and Theorem 1.2 in [DFD]. Corollary 1.2 can be regarded as a generalization of their results on Einstein manifolds [CGT, DF, Li3].
Based on some existing results in this field and Theorem 1.1, combined with Li’s conjecture that a closed $n$-dimensional Riemannian manifold with $(n+\frac{n-2}{n})$-positive curvature operator of the second kind is diffeomorphic to a spherical space form, it is natural to ask:
Question Are closed manifolds with harmonic curvature and $(n+\frac{n-2}{n})$-positive curvature operator of the second kind spherical space form?
2. proof of Theorem 1.1
In a Riemannian manifold $(M^{n},g)$, the Riemannian curvature tensor $R$ induces two symmetric linear operators $\hat{R}$ and $\mathring{R}$. Denote by $V$ be the tangent space of this Riemannian manifold, by ${\Lambda}^{2}V$ the space of skew symmetric 22-tensor over $V$, and by $S_{0}^{2}(V)$ the space of trace-free symmetric 22-tensor over $V$. The curvature operator of the first kind is defined by
$\hat{R}:{\Lambda}^{2}V\to{\Lambda}^{2}V$
$\hat{R}(e_{i}\wedge e_{j})=\frac{1}{2}\sum\limits_{k,l}{R_{ijkl}e_{k}\wedge e_{l}},$
and the curvature operator of the second kind is defined in [CGT, NP] by
$\mathring{R}={\Pr}_{s_{0}^{2}(V)}\circ\left.\bar{R}\right|_{S_{0}^{2}(V)}.$
The operator $\mathring{R}$ can be regarded as the following curvature operator $\bar{R}$ and its image restricts on $S_{0}^{2}(V)$. The operator $\bar{R}$ act on $S^{2}(V)$ induced from Riemannian curvature tensor $R$ is defined by
$\bar{R}(e_{i}\odot e_{j})=\sum\limits_{k,l}{R_{kijl}e_{k}\odot e_{l}},$
here $e_{i}\odot e_{j}=e_{i}\otimes e_{j}+e_{j}\otimes e_{i}$. So $\frac{1}{\sqrt{2}}\{e_{i}\odot e_{j}\}_{1\leq i<j\leq n},\frac{1}{2}\{e_{i}\odot e_{i}\}_{i=1,\dots,n}$ is an orthonormal basis for $S^{2}(V)$.
Let $M^{n}$ be an $n$-dimensional complete manifold with harmonic curvature, endowed with a standard Riemannian metric $\left\langle\text{,}\right\rangle$. Riemannian curvature tensor $R$ is divergence free if and only if its Ricci tensor is a Codazzi tensor, and in this case its scalar curvature is constant. Hence the traceless Ricci tensor is harmonic because its Ricci tensor is divergence free. Denote by $E$ its traceless Ricci tensor and by $s$ its scalar curvature. Then
$\Delta\frac{1}{2}\left|E\right|^{2}=\left|\nabla E\right|^{2}+\left\langle\Delta E,E\right\rangle.$ (2.1)
Since $E$ is harmonic, the Ricci identity gives
$\left\langle\Delta E,E\right\rangle$ $=E_{ij}E_{ij,tt}=E_{ij}E_{it,jt}$
$=E_{ij}(E_{it,tj}+R_{jtsi}E_{st}+R_{jtst}E_{is})$
$=\left\langle\mathring{R}\,(E),E\right\rangle+R_{ij}E_{it}E_{jt}.$ (2.2)
To analyze the individual terms, we apply the powerful weighted-sum calculus developed in [NP], which allows us to estimate finite weighted sums with nonnegative weights.
Definition 2.1.
([NP, Definition 3.1]) Let $\{\omega_{i}\}_{i=1}^{N}$ be the nonnegative weights of any finite weighted sums. Define
$\Omega=\max_{1\leq i\leq N}\omega_{i}\quad\mathrm{and}\quad S=\sum_{i=1}^{N}\omega_{i}\,.$
We call $S$ the total weight and $\Omega$ the highest weight.
Let $\{\lambda_{i}\}_{i=1}^{N}$ be the eigenvalues of $\mathring{R}\,$. Following the notation in [NP], we write $\mathring{R},\Omega,S]$ to denote any finite weighted sums $\sum_{i=1}^{N}\omega_{i}\lambda_{i}$ in terms of $\{\lambda_{i}\}_{i=1}^{N}$ whose weights satisfy highest weight $\Omega$ and total weight $S$.
Lemma $2.$
([NP, Lemma 3.4]) Let $\{\lambda_{i}\}_{i=1}^{N}$ denote the ordered eigenvalues of $\mathring{R}$. Then, for any integer $1\leq m\leq N$, we have
$\mathring{R},\Omega,S]\geq(S-m\Omega)\lambda_{m+1}+\Omega\sum_{i=1}^{m}\lambda_{i}.$
Therefore, $\mathring{R},\Omega,S]\geq 0$ if $\mathring{R}$ is $\frac{S}{\Omega}$-nonnegative.
If $\{S_{\alpha}\}$ is an orthonormal eigenbasis for $\mathring{R}$ with corresponding eigenvalues $\{\lambda_{\alpha}\}$, then since $E$ is a traceless symmetric $(0,2)$-tensor, it can be expressed in the orthonormal basis $\left\{S_{\alpha}\right\}$ as $E=\sum\nolimits_{\alpha}{E_{\alpha}S_{\alpha}}$. Then equation (2) becomes
$\left\langle\Delta E,E\right\rangle=R_{ij}E_{it}E_{jt}+\sum\limits_{\alpha}{{\lambda}_{\alpha}E_{\alpha}^{2}}.$ (2.3)
According to [NP, Example 3.2], the scalar curvature $s$ of $M$ satisfies
$s\geq\frac{2n}{n+2}[\mathring{R},1,\frac{(n-1)(n+2)}{2}],$ (2.4)
and from [NP, Lemma 3.14], the Ricci curvature RicRic of $M$ satisfies
$\operatorname{Ric}\geq\frac{n-1}{n+1}\left[\mathring{R},1,n\right]+\frac{1}{n(n+1)}s.$
Combining (2.4) with [NP, Lemma 3.3(c)], we obtain
$\operatorname{Ric}$ $\geq\frac{n-1}{n+1}\left[\mathring{R},1,n\right]+\frac{2}{(n+1)(n+2)}[\mathring{R},1,\frac{(n-1)(n+2)}{2}]$
$\geq\left[\mathring{R},\frac{n}{n+2},n-1\right].$ (2.5)
Furthermore, since
$\sum\limits_{\alpha}{E_{\alpha}^{2}}=\left|E\right|^{2}\quad\text{and}\quad E_{\alpha}^{2}\leq\left|E\right|^{2},$
the weight principle in [NP, Theorem 3.6] yields
$\left\langle\mathring{R}\,(E),E\right\rangle\geq\left[\mathring{R},1,1\right]\left|E\right|^{2}.$ (2.6)
Proposition 2.3.
If $E$ is a traceless symmetric 2-tensor, then
$\left\langle\Delta E,E\right\rangle\geq\left[\mathring{R},\frac{2(n+1)}{n+2},n\right]\left|E\right|^{2}.$
Proof.
Combining equation (2.3) with (2.5),(2.6) and [NP, Lemma 3.3(c)] yields
$\left\langle\Delta E,E\right\rangle\geq$ $\left[\mathring{R},\frac{n}{n+2},n-1\right]\left|E\right|^{2}+\left[\mathring{R},1,1\right]\left|E\right|^{2}$
$\geq$ $\left[\mathring{R},\frac{2(n+1)}{n+2},n\right]\left|E\right|^{2}.$
Proof of Theorem $1.$
[…]
$\left\langle\Delta E,E\right\rangle\geq\left[\mathring{R},\frac{2(n+1)}{n+2},n\right]\left|E\right|^{2}\geq 0.$
It follows from (2.1) that the trace-less Ricci tensor $E$ is parallel. Hence $\tilde{M}$ is Ricci parallel and an Einstein manifold. Consequently, $M$ is an Einstein manifold. ∎

### v2

@@
 Abstract.
-We prove that compact Riemannian manifolds of dimension $n\geq 3$ with harmonic curvature and $\frac{n(n+2)}{2(n+1)}$-nonnegative curvature operator of the second kind must be Einstein.
-In particular, Building upon Dai-Fu’s work [DF], it follows that if the curvature operator of the second kind is $\min\{\frac{n(n+2)}{2(n+1)},\max\{4,\frac{(n+2)}{4}\}\}$-nonnegative , then such a manifold must be of constant curvature.
+We prove that complete Riemannian manifolds of dimension $n\geq 3$ with harmonic curvature and $\frac{n(n+2)}{2(n+1)}$-nonnegative curvature operator of the second kind must be Einstein.
+In particular, We show that complete Einstein manifolds of dimension $n\geq 4$ with $\frac{3n(n-1)^{2}(n+2)}{2(5n^{3}+3n^{2}-30n+16)}$-nonnegative curvature operator of the second kind must be of constant curvature, which generalizes the work of Dai-Fu [DF24].
+Key words and phrases:
+Einstein manifolds, Bochner technique, Harmonic curvature
+2020 Mathematics Subject Classification:
+53C20, 53C24, 53C25.
 Supported in part by National Natural Science Foundations of China #12461008 and 12271069, Jiangxi Province Natural Science Foundation of China #20202ACB201001, Jiangxi Province Graduate Student Innovation Special Fund Project #YC2025-B037.
@@
 It is an important topic in geometry to understand how geometric assumptions restrict the topology of the underlying Riemannian manifold.
-In this direction, Tachibana [Ta] showed that any compact Riemannian manifold of dimension $n\geq 3$ with harmonic curvature and positive curvature operator is isometric to a quotient of the standard sphere.
+In this direction, Tachibana [Tac74] showed that any compact Riemannian manifold of dimension $n\geq 3$ with harmonic curvature and positive curvature operator is isometric to a quotient of the standard sphere.
 If the curvature operator is only nonnegative, then the manifold is locally symmetric.
-Later, Tran [tran] proved that a compact Riemannian manifold $M$ of dimension $n\geq 4$ with harmonic Weyl curvature and positive curvature operator, then $M$ is locally conformally flat.
-In [PW1] Petersen and Wink showed that a compact manifold with harmonic Weyl tensor and $\frac{n-1}{2}]-$nonnegative curvature operator is either globally conformal to a quotient of the standard sphere or locally symmetric, and that the first possibility always occurs if the curvature operator is $\frac{n-1}{2}]-$positive.
-In [PW] they proved a similar Tachibana-type theorem under the stronger condition that the manifold be Einstein.
-More recently, Colombo, Mariani, and Rigoli [CMR] established that a compact Riemannian manifold of dimension $n\geq 3$ with harmonic curvature and $\frac{n-1}{2}]-$positive curvature operator must have constant sectional curvature, thereby generalizing Tachibana’s result.
-For curvature operator of the second kind, Nishikawa [Ni] conjectured that a closed simply connected Riemannian manifold for which the curvature operator of the second kind is nonnegative and positive is diffeomorphic to a Riemannian locally symmetric space and a spherical space form, respectively.
-This conjecture has been proved and improved by Cao–Gursky–Tran [CGT] and Li [Li4], who relaxed the assumption to 33-positive (33-nonnegative) curvature operator of the second kind.
-In 1993, Kashiwada [ka] showed that Riemannian manifolds with harmonic curvature for which the curvature operator of the second kind is nonnegative (respectively, positive) are locally symmetric spaces (respectively, constant curvature spaces).
-Nienhaus, Petersen and Wink [NP, NPWW] used the Bochner formula to show that $n$-dimensional compact Einstein manifolds with $k(<\frac{3n(n+2)}{2(n+4)})$-nonnegative curvature operators of the second kind are either rational homology spheres or flat.
-Dai-Fu [DF] proved that any closed Einstein manifold of dimension $n=4$ or $n=5$ or $n\geq 8$ with $\frac{n+2}{4}]-$nonnegative curvature operator of the second kind is a constant curvature space.
-Later, Dai-Fu-Yang [DFD] proved a compact manifold with harmonic Weyl tensor and nonnegative curvature operator of the second kind is globally conformally equivalent to a space of positive constant curvature or is isometric to a flat manifold.
+Later, Tran [Tra17] proved that a compact Riemannian manifold $M$ of dimension $n\geq 4$ with harmonic Weyl curvature and positive curvature operator, then $M$ is locally conformally flat.
+In [PW22] Petersen and Wink showed that a compact manifold with harmonic Weyl tensor and $\frac{n-1}{2}]-$nonnegative curvature operator is either globally conformal to a quotient of the standard sphere or locally symmetric, and that the first possibility always occurs if the curvature operator is $\frac{n-1}{2}]-$positive.
+In [PW21] they proved a similar Tachibana-type theorem under the stronger condition that the manifold be Einstein.
+More recently, Colombo, Mariani, and Rigoli [CMR24] established that a compact Riemannian manifold of dimension $n\geq 3$ with harmonic curvature and $\frac{n-1}{2}]-$positive curvature operator must have constant sectional curvature, thereby generalizing Tachibana’s result.
+This result was improved by Bettiol and Goodman [BG24] to $\frac{n-1}{2}$-positive curvature operator for $n\geq 5$ and $\frac{n}{2}$-positive curvature operator for $n=3,4$.
+For curvature operator of the second kind, Nishikawa [Nis86] conjectured that a closed simply connected Riemannian manifold for which the curvature operator of the second kind is nonnegative and positive is diffeomorphic to a Riemannian locally symmetric space and a spherical space form, respectively.
+This conjecture has been proved and improved by Cao–Gursky–Tran [CGT23] and Li [Li24], who relaxed the assumption to $3$-positive (33-nonnegative) curvature operator of the second kind.
+In 1993, Kashiwada [Kas93] showed that Riemannian manifolds with harmonic curvature for which the curvature operator of the second kind is nonnegative (respectively, positive) are locally symmetric spaces (respectively, constant curvature spaces).
+Nienhaus, Petersen and Wink [NPW23, NPWW23] used the Bochner formula to show that $n$-dimensional compact Einstein manifolds with $k(<\frac{3n(n+2)}{2(n+4)})$-nonnegative curvature operators of the second kind are either rational homology spheres or flat.
+Dai-Fu [DF24] proved that any compact Einstein manifold must have constant curvature if its curvature operator of the second kind is $k$-nonnegative, where $k=1$ for $4\leq n\leq 7$, $k=2$ for $8\leq n\leq 10$ and $k=[\frac{n+2}{4}]$ for $n\geq 11$.
+Later, Dai-Fu-Yang [DFY24] proved a compact manifold with harmonic Weyl tensor and nonnegative curvature operator of the second kind is globally conformally equivalent to a space of positive constant curvature or is isometric to a flat manifold.
 Let $\lambda_{i}$ be the eigenvalues of the curvature operator, arranged in non-decreasing order. If
@@
 nonnegative).
-In this paper, we investigate Riemannian manifolds of dimension $n\geq 3$ with harmonic curvature and $\frac{n(n+2)}{2(n+1)}$-nonnegative curvature operator of the second kind, and obtain the following Theorem $1.$
-Theorem $1.$
-Let $(M,g)$ be an $n(\geq 3)$-dimensional compact Riemannian manifold with harmonic curvature.
+In this paper, we investigate Riemannian manifolds of dimension $n\geq 3$ with harmonic curvature and $\frac{n(n+2)}{2(n+1)}$-nonnegative curvature operator of the second kind, and obtain the following Theorem 1.1.
+Theorem 1.1.
+Let $(M,g)$ be an $n(\geq 3)$-dimensional complete Riemannian manifold with harmonic curvature tensor.
 If the curvature operator of the second kind $\mathring{R}$ is $\frac{n(n+2)}{2(n+1)}$-nonnegative, then $M$ is an Einstein manifold.
-Furthermore, by employing rigidity theorems for Einstein manifolds concerning the curvature operator of the second kind, established by Dai–Fu [DF] and Li [Li3], we get the following:
+For Einstein manifolds, we establish the following theorem, which generalizes the result of [DF24].
+Theorem 1.2.
+Let $(M,g)$ be an $n(\geq 4)$-dimensional complete Einstein manifold.
+If the curvature operator of the second kind $\mathring{R}$ is $\frac{3n(n-1)^{2}(n+2)}{2(5n^{3}+3n^{2}-30n+16)}$-nonnegative, then $M$ must be of constant curvature.
+Therefore, Based on Theorems 1.1 and  1.2, and the rigidity results for Einstein manifolds related to the curvature operator of the second kind — as established by Dai-Fu [DF24] and Li [Li22], we obtain the following corollary.
 Corollary 1.3.
-Let $(M,g)$ be an $n(\geq 3)$-dimensional compact Riemannian manifold with harmonic curvature, and $\mathring{R}\,$ be the curvature operator of the second kind. If
+Let $(M,g)$ be an $n(\geq 3)$-dimensional complete Riemannian manifold with harmonic curvature tensor, and $\mathring{R}\,$ be the curvature operator of the second kind. If
 (i) $3\leq n\leq 8$ and $\mathring{R}\,$ is $\frac{n(n+2)}{2(n+1)}$-nonnegative;
-(ii) $8\leq n\leq 13$ and $\mathring{R}\,$ is 44-nonnegative;
-(iii) $n\geq 14$, $\mathring{R}\,$ is $\frac{n+2}{4}]-$nonnegative,
+(ii) $9\leq n\leq 15$ and $\mathring{R}\,$ is $4\frac{1}{2}$-nonnegative;
+(iii) $n\geq 16$, $\mathring{R}\,$ is $\frac{3n(n-1)^{2}(n+2)}{2(5n^{3}+3n^{2}-30n+16)}$-nonnegative,
 then $M$ must be of constant curvature.
 Remark 1.4.
-Corollary 1.2 extends and enhances Kashiwada’s results [ka] and Theorem 1.2 in [DFD]. Corollary 1.2 can be regarded as a generalization of their results on Einstein manifolds [CGT, DF, Li3].
-Based on some existing results in this field and Theorem 1.1, combined with Li’s conjecture that a closed $n$-dimensional Riemannian manifold with $(n+\frac{n-2}{n})$-positive curvature operator of the second kind is diffeomorphic to a spherical space form, it is natural to ask:
+Corollary 1.3 extends and enhances Kashiwada’s results [Kas93] and Theorem 1.2 in [DFY24]. Corollary 1.3 can be regarded as a generalization of their results on Einstein manifolds [CGT23, DF24, Li22].
+Based on some existing results in this field and Corollary 1.3, combined with Li’s conjecture that a closed $n$-dimensional Riemannian manifold with $(n+\frac{n-2}{n})$-positive curvature operator of the second kind is diffeomorphic to a spherical space form, it is natural to ask:
 Question Are closed manifolds with harmonic curvature and $(n+\frac{n-2}{n})$-positive curvature operator of the second kind spherical space form?
-2. proof of Theorem 1.1
+2. notation and lemmas
 In a Riemannian manifold $(M^{n},g)$, the Riemannian curvature tensor $R$ induces two symmetric linear operators $\hat{R}$ and $\mathring{R}$.
-Denote by $V$ be the tangent space of this Riemannian manifold, by ${\Lambda}^{2}V$ the space of skew symmetric 22-tensor over $V$, and by $S_{0}^{2}(V)$ the space of trace-free symmetric 22-tensor over $V$.
+Denote by $V$ be the tangent space of this Riemannian manifold, by ${\Lambda}^{2}V$ the space of skew symmetric $2$-tensor over $V$, and by $S_{0}^{2}(V)$ the space of trace-free symmetric $2$-tensor over $V$.
 The curvature operator of the first kind is defined by
@@
 $\hat{R}(e_{i}\wedge e_{j})=\frac{1}{2}\sum\limits_{k,l}{R_{ijkl}e_{k}\wedge e_{l}},$
-and the curvature operator of the second kind is defined in [CGT, NP] by
+and the curvature operator of the second kind is defined in [CGT23, NPW23] by
 $\mathring{R}={\Pr}_{s_{0}^{2}(V)}\circ\left.\bar{R}\right|_{S_{0}^{2}(V)}.$
@@
 The operator $\bar{R}$ act on $S^{2}(V)$ induced from Riemannian curvature tensor $R$ is defined by
-$\bar{R}(e_{i}\odot e_{j})=\sum\limits_{k,l}{R_{kijl}e_{k}\odot e_{l}},$
+$\bar{R}(e_{i}\odot e_{j})=\sum\limits_{k,l}{R_{kijl}e_{k}\odot e_{l}}.$
 Here $e_{i}\odot e_{j}=e_{i}\otimes e_{j}+e_{j}\otimes e_{i}$.
 So $\frac{1}{\sqrt{2}}\{e_{i}\odot e_{j}\}_{1\leq i<j\leq n},\frac{1}{2}\{e_{i}\odot e_{i}\}_{i=1,\dots,n}$ is an orthonormal basis for $S^{2}(V)$.
+Definition 2.1.
+Let $T^{(0,k)}(V)$ denote the space of $(0,k)$-tensor space on $V$. For $S\in S^{2}(V)$ and $T\in T^{(0,k)}(V),$ we define
+$S:T^{(0,k)}(V)\to T^{(0,k)}(V)$
+$(ST)(X_{1},\cdots,X_{k})=\sum\limits_{i=1}^{k}{T(X_{1},\cdots,SX_{i},\cdots,X_{k})}$
+and define $T^{S^{2}}\in T^{(0,k)}(V)\otimes S^{2}(V)$ by
+$\left\langle T^{S^{2}}(X_{1},\cdots,X_{k}),S\right\rangle=(ST)(X_{1},\cdots,X_{k}).$
+Hence, if $\left\{S^{\alpha}\right\}$ is an orthonormal basis for $S_{0}^{2}(V)$, then
+$T^{S_{0}^{2}}=\sum\limits_{\alpha=1}^{N}{S^{\alpha}T\otimes S^{\alpha}}.$
+To analyze the individual terms for Bochner formula, we apply the powerful weighted-sum calculus developed in [NPW23], which allows us to estimate finite weighted sums with nonnegative weights.
+Definition 2.2.
+([NPW23, Definition 3.1]) Let $\{\omega_{i}\}_{i=1}^{N}$ be the nonnegative weights of any finite weighted sums. Define
+$\Omega=\max_{1\leq i\leq N}\omega_{i}\quad\mathrm{and}\quad S=\sum_{i=1}^{N}\omega_{i}\,.$
+We call $S$ the total weight and $\Omega$ the highest weight.
+Let $\{\lambda_{i}\}_{i=1}^{N}$ be the eigenvalues of $\mathring{R}\,$.
+Following the notation in [NPW23], we write $\mathring{R},\Omega,S]$ to denote any finite weighted sums $\sum_{i=1}^{N}\omega_{i}\lambda_{i}$ in terms of $\{\lambda_{i}\}_{i=1}^{N}$ whose weights satisfy highest weight $\Omega$ and total weight $S$.
+Lemma 2.3.
+([NPW23, Lemma 3.4]) Let $\{\lambda_{i}\}_{i=1}^{N}$ denote the ordered eigenvalues of $\mathring{R}$. Then, for any integer $1\leq m\leq N$, we have
+$\mathring{R},\Omega,S]\geq(S-m\Omega)\lambda_{m+1}+\Omega\sum_{i=1}^{m}\lambda_{i}.$
+Therefore, $\mathring{R},\Omega,S]\geq 0$ if $\mathring{R}$ is $\frac{S}{\Omega}$-nonnegative.
+3. manifold with harmonic curvature
 Let $M^{n}$ be an $n$-dimensional complete manifold with harmonic curvature, endowed with a standard Riemannian metric $\left\langle\text{,}\right\rangle$.
@@
 Then
-$\Delta\frac{1}{2}\left|E\right|^{2}=\left|\nabla E\right|^{2}+\left\langle\Delta E,E\right\rangle.$ (2.1)
+(3.1) $\frac{1}{2}\Delta\left|E\right|^{2}=\left|\nabla E\right|^{2}+\left\langle\Delta E,E\right\rangle.$
 Since $E$ is harmonic, the Ricci identity gives
@@
 $=E_{ij}(E_{it,tj}+R_{jtsi}E_{st}+R_{jtst}E_{is})$
-$=\left\langle\mathring{R}\,(E),E\right\rangle+R_{ij}E_{it}E_{jt}.$ (2.2)
-To analyze the individual terms, we apply the powerful weighted-sum calculus developed in [NP], which allows us to estimate finite weighted sums with nonnegative weights.
-Definition 2.1.
-([NP, Definition 3.1]) Let $\{\omega_{i}\}_{i=1}^{N}$ be the nonnegative weights of any finite weighted sums. Define
-$\Omega=\max_{1\leq i\leq N}\omega_{i}\quad\mathrm{and}\quad S=\sum_{i=1}^{N}\omega_{i}\,.$
-We call $S$ the total weight and $\Omega$ the highest weight.
-Let $\{\lambda_{i}\}_{i=1}^{N}$ be the eigenvalues of $\mathring{R}\,$.
-Following the notation in [NP], we write $\mathring{R},\Omega,S]$ to denote any finite weighted sums $\sum_{i=1}^{N}\omega_{i}\lambda_{i}$ in terms of $\{\lambda_{i}\}_{i=1}^{N}$ whose weights satisfy highest weight $\Omega$ and total weight $S$.
-Lemma $2.$
-([NP, Lemma 3.4]) Let $\{\lambda_{i}\}_{i=1}^{N}$ denote the ordered eigenvalues of $\mathring{R}$. Then, for any integer $1\leq m\leq N$, we have
-$\mathring{R},\Omega,S]\geq(S-m\Omega)\lambda_{m+1}+\Omega\sum_{i=1}^{m}\lambda_{i}.$
-Therefore, $\mathring{R},\Omega,S]\geq 0$ if $\mathring{R}$ is $\frac{S}{\Omega}$-nonnegative.
+(3.2) $=\left\langle\mathring{R}\,(E),E\right\rangle+R_{ij}E_{it}E_{jt},$
 If $\{S_{\alpha}\}$ is an orthonormal eigenbasis for $\mathring{R}$ with corresponding eigenvalues $\{\lambda_{\alpha}\}$, then since $E$ is a traceless symmetric $(0,2)$-tensor, it can be expressed in the orthonormal basis $\left\{S_{\alpha}\right\}$ as $E=\sum\nolimits_{\alpha}{E_{\alpha}S_{\alpha}}$.
 Then equation (3) becomes
-$\left\langle\Delta E,E\right\rangle=R_{ij}E_{it}E_{jt}+\sum\limits_{\alpha}{{\lambda}_{\alpha}E_{\alpha}^{2}}.$ (2.3)
-According to [NP, Example 3.2], the scalar curvature $s$ of $M$ satisfies
-$s\geq\frac{2n}{n+2}[\mathring{R},1,\frac{(n-1)(n+2)}{2}],$ (2.4)
-and from [NP, Lemma 3.14], the Ricci curvature RicRic of $M$ satisfies
+(3.3) $\left\langle\Delta E,E\right\rangle=R_{ij}E_{it}E_{jt}+\sum\limits_{\alpha}{{\lambda}_{\alpha}E_{\alpha}^{2}}.$
+According to [NPW23, Example 3.2], the scalar curvature $s$ of $M$ satisfies
+(3.4) $s\geq\frac{2n}{n+2}[\mathring{R},1,\frac{(n-1)(n+2)}{2}],$
+and from [NPW23, Lemma 3.14], the Ricci curvature $Ric$ of $M$ satisfies
 $\operatorname{Ric}\geq\frac{n-1}{n+1}\left[\mathring{R},1,n\right]+\frac{1}{n(n+1)}s.$
-Combining (2.4) with [NP, Lemma 3.3(c)], we obtain
+Combining (3.4) with [NPW23, Lemma 3.3(c)], we obtain
 $\operatorname{Ric}$ $\geq\frac{n-1}{n+1}\left[\mathring{R},1,n\right]+\frac{2}{(n+1)(n+2)}[\mathring{R},1,\frac{(n-1)(n+2)}{2}]$
-$\geq\left[\mathring{R},\frac{n}{n+2},n-1\right].$ (2.5)
+(3.5) $\geq\left[\mathring{R},\frac{n}{n+2},n-1\right].$
 Furthermore, since
 $\sum\limits_{\alpha}{E_{\alpha}^{2}}=\left|E\right|^{2}\quad\text{and}\quad E_{\alpha}^{2}\leq\left|E\right|^{2},$
-the weight principle in [NP, Theorem 3.6] yields
-$\left\langle\mathring{R}\,(E),E\right\rangle\geq\left[\mathring{R},1,1\right]\left|E\right|^{2}.$ (2.6)
+the weight principle in [NPW23, Theorem 3.6] yields
+(3.6) $\left\langle\mathring{R}\,(E),E\right\rangle\geq\left[\mathring{R},1,1\right]\left|E\right|^{2}.$
 Proposition 3.1.
@@
 Proof.
-Combining equation (2.3) with (2.5),(2.6) and [NP, Lemma 3.3(c)] yields
+Combining equation (3.3) with (3.5),(3.6) and [NPW23, Lemma 3.3(c)] yields
 $\left\langle\Delta E,E\right\rangle\geq$ $\left[\mathring{R},\frac{n}{n+2},n-1\right]\left|E\right|^{2}+\left[\mathring{R},1,1\right]\left|E\right|^{2}$
-$\geq$ $\left[\mathring{R},\frac{2(n+1)}{n+2},n\right]\left|E\right|^{2}.$
-Proof of Theorem $1.$
+$=$ $\left[\mathring{R},\frac{2(n+1)}{n+2},n\right]\left|E\right|^{2}.$
+Proof of Theorem 1.1.
 […]
@@
 It follows from (3.1) that the trace-less Ricci tensor $E$ is parallel. Hence $\tilde{M}$ is Ricci parallel and an Einstein manifold. Consequently, $M$ is an Einstein manifold. ∎
+4. Einstein manifold
+This section derives a refined Bochner formula for Einstein manifolds, leading to a proof of Theorem 1.2. We begin by establishing a Bochner formula for the curvature tensor.
+Proposition 4.1.
+Let $(M,g)$ be an $n$-dimensional Einstein manifold, then the curvature operator $R$ of $M$ satisfies
+(4.1) $3\left\langle\Delta R,R\right\rangle=\sum\limits_{\alpha}{{\lambda}_{\alpha}\left|S^{\alpha}W\right|^{2}}+16\sum\limits_{\alpha}{{\lambda}_{\alpha}({\lambda}_{\alpha}-\frac{s}{n(n-1)})^{2}}+\frac{4(n-4)}{n^{2}}s\left|W\right|^{2}.$
+Proof.
+The following Bochner formula is established by Dai and Fu in [DF24, Proposition 3.4].
+$3\left\langle\Delta R,R\right\rangle=$ $\sum\limits_{\alpha}{{\lambda}_{\alpha}\left|S^{\alpha}W\right|^{2}}+8\left(\frac{-n^{3}+6n^{2}+12n-8}{3n^{4}(n-1)^{2}}\right)s^{3}$
+$+8\left(\frac{2n^{2}-22n+8}{3n^{2}(n-1)}\right)s\sum\limits_{\alpha}{\lambda_{\alpha}^{2}}+16\sum\limits_{\alpha}{\lambda_{\alpha}^{3}}.$
+From [DF24, Lemma 3.3], we know
+$tr(\mathring{R})$ $=$ $\frac{n+2}{2n}s,$
+(4.2) $tr(\mathring{R}^{2})$ $=$ $\frac{3}{4}|R|^{2}-\frac{s^{2}}{n^{2}}=\frac{3}{4}|W|^{2}+\frac{n+2}{2n^{2}(n-1)}s^{2}.$
+Combining the Bochner formula above with (4), we obtain
+$3\left\langle\Delta R,R\right\rangle=\sum\limits_{\alpha}{{\lambda}_{\alpha}\left|S^{\alpha}W\right|^{2}}+16\sum\limits_{\alpha}{{\lambda}_{\alpha}({\lambda}_{\alpha}-\frac{s}{n(n-1)})^{2}}+\frac{4(n-4)}{n^{2}}s\left|W\right|^{2}.$
+First, let’s estimate the first term on the right-hand side of (4.1). According to the calculations in [DF24], we have
+$\sum\limits_{\alpha=1}^{N}{\left|S^{\alpha}W\right|^{2}}=\frac{2(n^{2}+n-8)}{n}{\left|W\right|^{2}},$
+and
+$\left|S^{\alpha}W\right|^{2}\leq\frac{8(n-2)}{n}\left|W\right|^{2}.$
+Hence
+(4.3) $\sum\limits_{\alpha}{{\lambda}_{\alpha}\left|S^{\alpha}W\right|^{2}}\geq[\mathring{R},\frac{8(n-2)}{n},\frac{2(n^{2}+n-8)}{n}]|W|^{2}.$
+Second, we estimate the second term on the right-hand side of (4.1). Set $\beta_{\alpha}={\lambda}_{\alpha}-\frac{s}{n(n-1)}$. From (4), we obtain
+[…]
+For any fixed $k$, the Cauchy–Schwarz inequality give
+$(\sum_{\alpha\neq k}\beta_{\alpha})^{2}\leq(N-1)\sum_{\alpha\neq k}{\beta_{\alpha}^{2}},$
+and hence
+$\beta_{k}^{2}\leq(N-1)[\frac{3}{4}|W|^{2}-\beta_{k}^{2}].$
+This implies that
+$\beta_{k}^{2}\leq\frac{3}{4}\frac{N-1}{N}|W|^{2}.$
+Therefore, we obtain
+(4.4) $\sum\limits_{i}{{\lambda}_{i}({\lambda}_{i}-\frac{s}{n(n-1)})^{2}}\geq\frac{3}{4}[\mathring{R},\frac{n^{2}+n-4}{(n-1)(n+2)},1]|W|^{2}.$
+Proposition 4.2.
+Let $M$ be an $n$-dimensional Einstein manifold with $n\geq 4$. Then the curvature operator $R$ of $M$ satisfies
+$3\left\langle\Delta R,R\right\rangle\geq\left[\mathring{R},\frac{4(5n^{3}+3n^{2}-30n+16)}{n(n-1)(n+2)},6(n-1)\right]\left|W\right|^{2}.$
+Proof.
+Combining equation (4.1) with (3.4),(4.3),(4.4) and [NPW23, Lemma 3.3(c)] yields
+$3\left\langle\Delta R,R\right\rangle=\left\langle\mathring{R}\,(W^{S_{0}^{2}}),W^{S_{0}^{2}}\right\rangle+16\sum\limits_{i}{{\lambda}_{i}({\lambda}_{i}-\frac{s}{n(n-1)})^{2}}+\frac{4(n-4)}{n^{2}}s\left|W\right|^{2}$
+$\geq[\mathring{R},\frac{8(n-2)}{n},\frac{2(n^{2}+n-8)}{n}]\left|W\right|^{2}+12[\mathring{R},\frac{n^{2}+n-4}{(n-1)(n+2)},1]\left|W\right|^{2}+\frac{8(n-4)}{n(n+2)}[\mathring{R},1,\frac{(n-1)(n+2)}{2}]\left|W\right|^{2}$
+$\geq[\mathring{R},\frac{8(n-2)}{n}+\frac{n^{2}+n-4}{(n-1)(n+2)}+\frac{8(n-4)}{n(n+2)},\frac{2(n^{2}+n-8)}{n}+12+\frac{4(n-4)(n-1)}{n}]\left|W\right|^{2}$
+$=[\mathring{R},\frac{4(5n^{3}+3n^{2}-30n+16)}{n(n-1)(n+2)},6(n-1)]\left|W\right|^{2}.$
+Proof of Theorem 1.2..
+If $s=0$, then $M$ is flat since $\mathring{R}$ is $\frac{3n(n-1)^{2}(n+2)}{2(5n^{3}+3n^{2}-30n+16)}$-nonnegative. Otherwise, $s>0$, and $M$ is compact by Myers Theorem for $M$ is Einstein.
+If $\mathring{R}$ is $\frac{3n(n-1)^{2}(n+2)}{2(5n^{3}+3n^{2}-30n+16)}$-nonnegative, then by using Proposition 4.2 and Lemma 2.3, we get
+$\frac{1}{2}\Delta\left|R\right|^{2}=\left|\nabla R\right|^{2}+\left\langle\Delta R,R\right\rangle\geq 0.$
+By the maximum principle, we have $\nabla R=0$. It then follows that $M$ is a locally symmetric space.
+On the other hand, It follows from Theorem B of [NPW23] that M is a rational homology sphere.
+Wolf [Wol69, Theorem 1] provides a complete classification of compact symmetric spaces that are rational homology spheres.
+Aside from spheres themselves, the only simply connected example is the space $SU(3)/SO(3)$, which is shown in [NPW23, Example 4.5] to be 9-positive.
+It then follows that the manifold must have constant curvature.
+This completes the proof of Theorem 1.2.
+∎